OpenStudy (anonymous):

is this correct ? Factor: x2 - 13x + 36 (x - 4)(x - 9)

OpenStudy (turingtest):

yep :)

Directrix (directrix):

Yes. (x - 4)(x - 9) = x^2 - 4x - 9x + 36 = x^2 - 13x + 36 which is the original expression.